OpenStudy (anonymous):

Find the matrix A of the linear transformation T from R^2 to R^2 that rotates any vector through an angle of 30 degrees in the counterclockwise direction.

OpenStudy (anonymous):

\[A=\left[\begin{matrix}? & ? \\ ? & ?\end{matrix}\right]\]

OpenStudy (anonymous):

If you want to derive the \(2 \times 2\) rotation matrix by hand, think of how to transform each of the standard basis vectors.

OpenStudy (anonymous):

\(T(e_1) = T((1, 0)) = (\cos(\theta), \sin(\theta))\) \(T(e_2) = T((0, 1)) = (-\sin(\theta), \cos(\theta))\) Therefore the matrix is: \(\left(\begin{array}{cc} \cos(\theta) & -\sin(\theta) \\ \sin(\theta) & cos(\theta) \end{array}\right)\) Now set \(\theta\) to 30 degrees and you have your matrix.

OpenStudy (anonymous):

So you are sayign that my answer should be \[\left[\begin{matrix}\cos(30) & -\sin(30) \\ \sin(30) & \cos(30)\end{matrix}\right]\ because this is what I had as an answer but it says it is incorrect.

OpenStudy (anonymous):

\[\left[\begin{matrix}\cos(30) & -\sin(30) \\ \sin(30) & \cos(30)\end{matrix}\right]\]

OpenStudy (anonymous):

Evaluate \(\sin\) and \(\cos\) on 30 to get numeric answers.

OpenStudy (anonymous):

\(\sin(30^\circ) = 0.5\) \(\cos(30^\circ) = \frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

Try numeric values and see if it accepts that.

OpenStudy (anonymous):

It did. Thank you for your help.